# Class-Proportional Coreset Selection for Difficulty-Separable Data

### v1

@@ Abstract
High-quality training data is essential for building reliable and efficient machine learning systems. One-shot coreset selection addresses this by pruning the dataset while maintaining or even improving model performance, often relying on training-dynamics-based data difficulty scores. However, most existing methods implicitly assume class-wise homogeneity in data difficulty, overlooking variation in data difficulty across different classes. In this work, we challenge this assumption by showing that, in domains such as network intrusion detection and medical imaging, data difficulty often clusters by class. We formalize this as class-difficulty separability and introduce the Class Difficulty Separability Coefficient (CDSC) as a quantitative measure. We demonstrate that high CDSC values correlate with performance degradation in class-agnostic coreset methods, which tend to overrepresent easy majority classes while neglecting rare but informative ones. To address this, we introduce class-proportional variants of multiple sampling strategies. Evaluated on five diverse datasets spanning security and medical domains, our methods consistently achieve state-of-the-art data efficiency. For instance, on CTU-13, at an extreme $99\%$ pruning rate, a class-proportional variant of Coverage-centric Coreset Selection (CCS-CP) shows remarkable stability, with accuracy dropping only $2.58\%$, precision $0.49\%$, and recall $0.19\%$. In contrast, the class-agnostic CCS baseline, the next best method, suffers sharper declines of $7.59\%$ in accuracy, $4.57\%$ in precision, and $4.11\%$ in recall. We further show that aggressive pruning enhances generalization in noisy, imbalanced, and large-scale datasets. Our results underscore that explicitly modeling class-difficulty separability leads to more effective, robust, and generalizable data pruning, particularly in high-stakes scenarios.
@@ 1 Introduction
Coreset selection and data pruning have emerged as effective tools for reducing the size of training datasets while preserving—or even improving—model performance. This paradigm has seen success across image tasks such as classification and object detection [28, 23, 5, 38, 37, 21, 33, 3, 40, 17, 15], as well as recent applications in large language model (LLM) pretraining and supervised fine-tuning [32, 31, 2, 19, 39, 14]. However, despite this growing body of work, the underlying assumptions behind many coreset selection methods remain largely unexamined.
@@
To characterize the extent of difficulty variation across classes, we introduce the Class Difficulty Separability Coefficient (CDSC), a normalized Jensen–Shannon divergence [18] computed over class-conditional data difficulty distributions. CDSC quantifies how distinctly different the difficulty scores are across classes. In Section 5.2.1, we show that class-agnostic coreset selection methods degrade more significantly on datasets with higher $\delta_{\text{CD}}$, particularly in terms of accuracy, precision, and recall. This supports our hypothesis that modeling class-specific difficulty becomes increasingly important as difficulty separability increases.
In this work, we investigate how class-difficulty separability impacts the effectiveness of one-shot coreset selection and propose a simple yet effective remedy: class-proportional variants of existing sampling strategies. Specifically, we adapt several data difficulty–based methods—including hardest example [24, 12], sliding-window [3, 40], and Coverage-Centric Coreset Selection (CCS) [38]—by applying coverage within each class and allocating selection quotas proportionally to the number of examples in each class. We show that this modification consistently improves the performance of each method. Our approach is modular and can be seamlessly integrated with other training-dynamics-based selectors.
@@
  • We identify and formalize class-difficulty separability as a structural property that can degrade the performance of class-agnostic coreset selection methods. To quantify it, we introduce the Class Difficulty Separability Coefficient (CDSC), a normalized Jensen–Shannon divergence over class-wise difficulty distributions. Higher CDSC values indicate greater performance gaps between class-agnostic and class-aware methods (Sec 3.35.2.1).
@@ 2.1 Coreset Selection
Coreset selection aims to find a small subset of training data that retains the performance of models trained on the full dataset. Existing methods fall into two main categories: (1) training-free methods, such as geometry or diversity-based selection [1, 30, 27, 33]; and (2) training-dynamics–based methods [24, 28, 23], which use training behavior to estimate data importance. We focus on the latter due to their superior performance.
@@ 2.2 Difficulty-Separable Data
Many real-world datasets exhibit inter-class difficulty divergence—some classes are consistently easier to learn, while others are inherently harder due to factors such as greater intra-class variability, ambiguous boundaries, or lower representation. This challenges standard, class-agnostic coreset selection methods. We highlight two domains where this structure is especially prominent:
@@ 21      $\mathcal{S}_{c}\leftarrow\mathcal{S}_{c}\cup S_{c}$ ;
where $H(P)=-\sum_{j}P_{j}\log P_{j}$ is the Shannon entropy of distribution $P$. The coefficient $\delta_{\text{CD}}$ takes the value 00 when all class-conditional difficulty distributions are identical, and approaches $1$ as their supports become mutually disjoint.
@@ 4.2 Baselines and Evaluation Metrics
We compare our class-proportional variants against several widely used coreset selection strategies: 1) Random samples data uniformly at random and serves as a naive, class-agnostic baseline; 2) Hardest [24, 28] selects the most difficult examples according to a training-dynamics–based difficulty score (e.g., AUM or forgetting score); 3) Sliding Window (SW) [3, 40] selects examples from an interval $[\beta,\beta+\alpha]$ of the sorted difficulty scores, where $\alpha$ is the target pruning rate and $\beta$ is a tunable starting offset; 4) Coverage-Centric Sampling (CCS) [38] discards the top $\beta\%$ hardest examples and then applies stratified sampling over the remaining data to select an $\alpha$-fraction coreset.
@@ 5.1 Main Evaluation
On security datasets with high class-difficulty separability (see Section 5.2.1), CCS-CP consistently delivers the strongest performance across accuracy, precision, and recall—especially under aggressive pruning. On CTU-13, CCS-CP matches the full-data accuracy at 90% pruning and even improves recall ($+0.53\%$) while keeping the precision drop minimal ($-0.07\%$). At an extreme $99\%$ pruning rate, CCS-CP maintains remarkable stability, with only a $2.58\%$ drop in accuracy, $0.49\%$ in precision, and $0.19\%$ in recall. In contrast, the class-agnostic CCS baseline—the next best method in terms of overall accuracy, precision, and recall—suffers much steeper declines: $7.59\%$ in accuracy, $4.57\%$ in precision, and $4.11\%$ in recall. Similar patterns hold across UNSW-NB15, CICIDS2017, DermaMNIST, and BloodMNIST, where CCS-CP maintains stronger accuracy, precision, and recall even under high pruning rates. This underscores the strength of class-proportional selection in preserving high-quality, representative examples from all classes. In contrast, class-agnostic methods may inadvertently discard critical minority classes, leading to degraded performance, particularly harmful for imbalanced datasets common in medical image classification and security domains such as network intrusion detection.
@@ Appendix B Implementation
[…]
@@ B.1 Hierarchical Beta Search
The search is conducted in three phases, each with a finer step size: $\delta=0.1$, $0.05$, and $0.01$. These granularities are chosen based on the smallest coreset ratio used in most datasets, which is $0.01$. Each phase begins at the $\beta$ found in the previous phase ($\beta=0$ for phase 00) and moves forward by incrementally adding $\delta$. The algorithm continues as long as validation accuracy improves. Once a decline is observed, it backtracks by $1.5\times\delta$ and proceeds to the next phase. To reduce redundant computation, we cache performance results for previously evaluated $\beta$ values. The search terminates after completing all three phases and returns the $\beta$ value that yielded the best validation accuracy.

### v2

@@ Abstract
High-quality training data is essential for building reliable and efficient machine learning systems. One-shot coreset selection addresses this by pruning the dataset while maintaining or even improving model performance, often relying on training-dynamics-based data difficulty scores. However, most existing methods implicitly assume class-wise homogeneity in data difficulty, overlooking variation in data difficulty across different classes. In this work, we challenge this assumption by showing that, in domains such as network intrusion detection and medical imaging, data difficulty often clusters by class. We formalize this as class-difficulty separability and introduce the Class Difficulty Separability Coefficient (CDSC) as a quantitative measure. We demonstrate that high CDSC values correlate with performance degradation in class-agnostic coreset methods, which tend to overrepresent easy majority classes while neglecting rare but informative ones. To address this, we introduce class-proportional variants of multiple sampling strategies. Evaluated on five diverse datasets spanning security and medical domains, our methods consistently achieve state-of-the-art performance. For instance, on CTU-13, at an extreme $99\%$ pruning rate, a class-proportional variant of Coverage-centric Coreset Selection (CCS-CP) shows remarkable stability, with accuracy dropping only $2.58\%$, precision $0.49\%$, and recall $0.19\%$. In contrast, the class-agnostic CCS baseline, the next best method, suffers sharper declines of $7.59\%$ in accuracy, $4.57\%$ in precision, and $4.11\%$ in recall. We further show that aggressive pruning enhances generalization in noisy, imbalanced, and large-scale datasets. Our results underscore that explicitly modeling class-difficulty separability leads to more effective, robust, and generalizable data pruning, particularly in high-stakes scenarios such as security and medical imaging.
@@ 1 Introduction
Coreset selection and data pruning have emerged as effective tools for reducing the size of training datasets while preserving, and in some cases even improving, model performance. This paradigm has seen success across vision tasks such as image classification and object detection [29, 23, 5, 42, 41, 21, 37, 3, 44, 17, 15], as well as recent applications in large language model (LLM) pretraining and supervised fine-tuning [36, 35, 2, 19, 43, 14]. However, despite this growing body of work, the underlying assumptions behind many coreset selection methods remain largely unexamined.
@@
To characterize the extent of difficulty variation across classes, we introduce the Class Difficulty Separability Coefficient (CDSC, Section 3.3), a normalized Jensen–Shannon divergence [18] computed over class-conditional data difficulty distributions. CDSC quantifies how distinctly different the difficulty scores are across classes. In Section 5.2.1, we show that class-agnostic coreset selection methods degrade more significantly on datasets with higher $\delta_{\text{CD}}$. This supports our hypothesis that modeling class-specific difficulty becomes increasingly important as difficulty separability increases.
In this work, we investigate how class-difficulty separability impacts the effectiveness of one-shot coreset selection and propose a simple yet effective remedy: class-proportional variants of existing sampling strategies. Specifically, we adapt several data difficulty–based methods, including hardest example [24, 12], sliding-window [3, 44], and Coverage-centric Coreset Selection (CCS) [42], by applying coverage within each class and allocating selection quotas proportionally to the number of examples in each class. We show that this modification consistently improves the performance of each method. Our approach is modular and can be seamlessly integrated with other training-dynamics-based selectors.
@@
  • We identify and formalize class-difficulty separability as a structural property that can degrade the performance of class-agnostic coreset selection methods. To quantify it, we introduce the Class Difficulty Separability Coefficient (CDSC), a normalized Jensen–Shannon divergence over class-wise difficulty distributions. Higher CDSC values indicate greater performance gaps between class-agnostic and class-aware methods.
@@ 2.1 Coreset Selection
Given a fixed budget, coreset selection seeks an optimal subset of training data that maximizes model performance, which can match or even surpass the performance achieved using the full dataset. Existing methods fall into two main categories: (1) training-free methods, such as geometry or diversity-based selection [1, 34, 28, 37]; and (2) training-dynamics–based methods [24, 29, 23], which use training behavior to estimate data importance. We focus on the latter due to its better empirical performance.
@@ 2.2 Difficulty-Separable Data
Machine learning has become a key component in addressing a wide range of security challenges [32, 25, 31, 30]. Many real-world datasets exhibit inter-class difficulty divergence—some classes are consistently easier to learn, while others are inherently harder due to factors such as greater intra-class variability, ambiguous boundaries, or lower representation. This challenges standard, class-agnostic coreset selection methods. We highlight two domains where this structure is especially prominent:
@@ 21 $\mathcal{S}_{c}\leftarrow\mathcal{S}_{c}\cup S_{c}$ ;
where $H(P)=-\sum_{j}P_{j}\log P_{j}$ is the Shannon entropy of distribution $P$. The coefficient $\delta_{\text{CD}}$ takes the value 0 when all class-conditional difficulty distributions are identical, and approaches $1$ as their supports become mutually disjoint.
@@ 4.2 Baselines and Evaluation Metrics
We compare our class-proportional variants against several widely used coreset selection strategies: 1) Random samples data uniformly at random and serves as a naive, class-agnostic baseline; 2) Hardest [24, 29] selects the most difficult examples according to a training-dynamics–based difficulty score (e.g., AUM or forgetting score); 3) Sliding Window (SW) [3, 44] selects examples from an interval $\beta,\beta+\alpha]$ of the sorted difficulty scores, where $\alpha$ is the target pruning rate and $\beta$ is a tunable starting offset; 4) Coverage-Centric Sampling (CCS) [42] discards the top $\beta\%$ hardest examples and then applies stratified sampling over the remaining data to select an $\alpha$-fraction coreset.
@@ 5.1 Main Evaluation
On security datasets with high class-difficulty separability (see Section 5.2.1), CCS-CP consistently delivers the strongest performance across accuracy, precision, and recall—especially under aggressive pruning. On CTU-13, CCS-CP matches the full-data accuracy at 90% pruning and even improves recall ($0.53\%$) while keeping the precision drop minimal ($0.07\%$). At an extreme $99\%$ pruning rate, CCS-CP maintains remarkable stability, with only a $2.58\%$ drop in accuracy, $0.49\%$ in precision, and $0.19\%$ in recall. In contrast, the class-agnostic CCS baseline—the next best method in terms of overall accuracy, precision, and recall—suffers much steeper declines: $7.59\%$ in accuracy, $4.57\%$ in precision, and $4.11\%$ in recall. Similar patterns hold across UNSW-NB15, CICIDS2017, DermaMNIST, and BloodMNIST, where CCS-CP maintains stronger accuracy, precision, and recall even under high pruning rates. This underscores the strength of class-proportional selection in preserving high-quality, representative examples from all classes. In contrast, class-agnostic methods may inadvertently discard critical minority classes, leading to degraded performance, particularly harmful for imbalanced datasets common in medical image classification and security domains such as network intrusion detection.
@@ Appendix B Implementation
All models are trained using the Adam optimizer [20] with early stopping based on validation accuracy. We scale the maximum number of training epochs inversely with the coreset sampling fraction to ensure fair comparisons across pruning levels. Learning rates are set to $2\text{e-}4$ for CTU-13 [9] and $1\text{e-}3$ for the rest. Batch sizes are 64 for UNSW-NB15 [22] and CTU-13 [9], 128 for DermaMNIST and BloodMNIST [38], and 512 for CICIDS2017 [27].
@@ B.1 Hierarchical Beta Search
The search is conducted in three phases, each with a finer step size: $\delta=0.1$, $0.05$, and $0.01$. These granularities are chosen based on the smallest coreset ratio used in most datasets, which is $0.01$. Each phase begins at the $\beta$ found in the previous phase ($\beta=0$ for phase 0) and moves forward by incrementally adding $\delta$. The algorithm continues as long as validation accuracy improves. Once a decline is observed, it backtracks by $1.5\times\delta$ and proceeds to the next phase. To reduce redundant computation, we cache performance results for previously evaluated $\beta$ values. The search terminates after completing all three phases and returns the $\beta$ value that yielded the best validation accuracy.